### v1

# Some stability results for the fractional differential equations with two delays

Abstract
This paper investigates the stability properties of a nonlinear fractional differential equation with two discrete delays and a delay-dependent coefficient. Such equations arise in various biological and control systems where temporal delays influence feedback mechanisms. Two cases are analyzed: one where the first delay is zero, and the second acts through an exponential coefficient, and another where the first delay is fixed. We derive delay-independent stability conditions using linearization, characteristic equations, and bifurcation theory, along with complete theoretical proofs. The results are validated through numerical simulations and stability diagrams.
1 Introduction
In 1695, Leibnitz proposed a notation $\frac{d^{n}x}{dt^{n}}$ for the $n^{th}$ order derivative of a function $x(t)$, where $n$ is a natural number. In his letter dated 30 September 1695, L’Hopital leibniz1860leibnizens ; podlubny1998fractional asked the meaning of this term when $n=1/2$. This is popularly known as the birth of the fractional calculus, a branch of mathematics that studies operators of noninteger order. By Fractional Derivative (FD), one means the derivative of an arbitrary order, e.g, a real or complex number, a function of t, etc.. The fractional derivative operators are non-local, unlike the classic integer-order derivative bhalekar2019can ; li2007remarks . Thus, one must provide the state values from the initial point when evaluating the FDs. This procedure proved extremely useful while modeling the memory properties in natural systems Podlubny1999 ; kilbas2006theory . This also gives the flexibility of selecting an order that fits the experimental data mainardi2022fractional ; bagley1983theoretical .
Similarly, the Delay Differential Equations (DDE) contain the states at the past time values smith2010introduction ; lakshmanan2011dynamics . Though the DDEs look very similar to their non-delayed counterparts, they are extremely difficult to analyze analytically hale2013introduction . A single scalar DDE can have oscillatory solutions. Moreover, one can have a scalar DDE exhibiting chaos sprott2007simple ; senthilkumar2005bifurcations ; ruiz2013chaos . The DDEs are infinite-dimensional dynamical systems that occur in various models in physics otto2019nonlinear ; bazighifan2021oscillation , chemistry epstein1991differential ; roussel1996use , biology and medicine rihan2021delay ; bani2017analysis ; glass2021nonlinear ; dong2023application , engineering and control systems insperger2011semi ; michiels2010control ; gumussoy2014computer , finance nagarajan2025stock ; flore2019feynman ; agrawal2020jump , and so on.
The fractional order differential equations with delay(FDDE) contain both these tools: the fractional derivative as well as the delay. Thus, one gets a better model for the given natural system. The numerical methods to solve FDDEs are presented in bhalekar2011predictor ; daftardar2014new . The stability and bifurcation analysis of these systems is developed in bhalekar2016stability ; bhalekar2011fractional ; gupta2024fractional ; bhalekar2022stability . The examples of chaotic FDDEs are bhalekar2012dynamical ; daftardar2012dynamics ; bhalekar2012generalized . The DDEs involving multiple delays occur in applications involving two or more processes gu2003stability ; belair1995age ; niculescu2002delay . The logistic model with two delays is studied in braddock1983two . Piotrowska proposed a model with two delays in tumor growth piotrowska2008hopf . Boullu et al employed such equations in the production of platelets boullu2020stability .
There are very few papers devoted to the analysis of such equations. Hale and Huang hale1993global , Belair and Campbell belair1994stability , Li, Ruan, and Weili1999stability , Bhalekar bhalekar2019analysing studied some particular cases of DDEs with multiple delays. Bhalekar and Dutta bhalekar2025analysis discussed the stability of a class of FDDEs with two delays. To the best of our knowledge, such equations with fractional derivatives have not been studied much in the literature. This motivates us to work on their stability and bifurcation analysis. In this work, we provide some delay-independent results.
This paper is organized as follows. Section 2 reviews preliminary results on fractional delay differential equations with a single discrete delay. In Section 3, we introduce the nonlinear fractional delay differential equation with delay-dependent coefficients, establish its equilibrium, and derive the corresponding linearized and characteristic equations. Sections 4 and 5 present the main stability theorems for the cases $\tau_{1}=0$ and $\tau_{1}=1$, together with illustrative examples and graphical results. Finally, Section 6 provides concluding remarks and outlines possible directions for future research.
2 Preliminaries
We consider a linear delay differential equation with Caputo fractional derivative $D^{\alpha}$ podlubny1998fractional ; kilbas2006theory .
Theorem 2.1.
bhalekar2016stability Consider the scalar FDE with a single discrete delay:
$D^{\alpha}x(t)=ax(t)+bx(t-\tau),\quad 0<\alpha\leq 1,\tau\geq 0.$ (1)
Then, the zero equilibrium $x_{*}=0$ has the following stability behavior:
  1. 1.
    If $b\in(-\infty,-|a|)$, then the equilibrium is asymptotically stable for $\tau\in[0,\tau_{cr})$ and the system undergoes Hopf bifurcation at
    $\tau_{cr}=\frac{\arccos\left(\frac{(a\cos(\frac{\alpha\pi}{2})+\sqrt{b^{2}-a^{2}\sin^{2}(\frac{\alpha\pi}{2})})\cos(\frac{\alpha\pi}{2})-a}{b}\right)}{\left(a\cos\left(\frac{\alpha\pi}{2}\right)+\sqrt{b^{2}-a^{2}\sin^{2}\left(\frac{\alpha\pi}{2}\right)}\right)^{1/\alpha}}.$
  2. 2.
    If $b\in(-a,\infty)$, then the equilibrium is unstable for all $\tau\geq 0$.
  3. 3.
    If $a<0$ and $b\in(a,-a)$, then the equilibrium is asymptotically stable for all $\tau\geq 0$.
Figure 1: Stability regions of Equation (1).
Notation: Here, SSR means Single Stable Region.
3 The Main Model
We consider the nonlinear fractional delay differential equation of the form:
$D^{\alpha}x(t)=-\gamma x(t)+g(x(t-\tau_{1}))-e^{-\gamma\tau_{2}}g(x(t-\tau_{1}-\tau_{2})),\quad 0<\alpha\leq 1,$ (2)
where $D^{\alpha}$ denotes the Caputo fractional derivative, $\gamma\in\mathbb{R}$, and $g\in C^{1}(\mathbb{R})$. The parameters $\tau_{1},\tau_{2}\geq 0$ represent discrete delays, and the exponential factor $e^{-\gamma\tau_{2}}$ introduces a delay-dependent coefficient. The integer-order counterpart of this equation, analyzed in boullu2020stability , represents a model for platelet production.
Let the equilibrium point of eq.(2) be $x_{*}$. Then, $x_{*}$ will be obtained by solving
$\gamma x_{*}+g(x_{*})-g(x_{*})e^{-\gamma\tau_{2}}=0.$
Here, we assume $g$ to be a nonlinear function such that $g(0)=0$ so that $x_{*}=0$ is an equilibrium point of eq.(2). Hence, assuming $g^{\prime}(0)=k$, we linearize equation (2) about the equilibrium $x_{*}=0$.
  1. 1.
    Linearization near Equilibrium
    Let $x(t)$ be an infinitesimally perturbed solution of (2) near $x_{*}=0.$ Using the first-order Taylor approximation for the nonlinear terms about the equilibrium and denoting $g^{\prime}(0)=k$, we have
    $g(x(t-\tau_{1}))\approx g(0)+g^{\prime}(0)x(t-\tau_{1})=kx(t-\tau_{1}),$
    $g(x(t-\tau_{1}-\tau_{2}))\approx g(0)+g^{\prime}(0)x(t-\tau_{1}-\tau_{2})=kx(t-\tau_{1}-\tau_{2}).$
    Substituting into equation (2), we obtain the linearized equation i.e.
    $D^{\alpha}x(t)=-\gamma x(t)+kx(t-\tau_{1})-ke^{-\gamma\tau_{2}}x(t-\tau_{1}-\tau_{2}).$ (3)
    We analyze the stability of this linearized equation in the presiding sections for various particular cases of parameters.
  2. 2.
    Characteristic Equation
    To analyze stability, we apply the Laplace transform and obtain the characteristic equation corresponding to equation (3) as
    $\lambda^{\alpha}=-\gamma+ke^{-\lambda\tau_{1}}-ke^{-\gamma\tau_{2}}e^{-\lambda(\tau_{1}+\tau_{2})}.$ (4)
We investigate the stability properties of the linearized system (3) in the following particular cases:
  • Case 1: $\tau_{1}=0$ and arbitrary $\tau_{2}>0.$
  • Case 2: $\tau_{1}=1$ and arbitrary $\tau_{2}>0.$
4 Case 1: $\tau_{1}=0$
Substituting $\tau_{1}=0$, $\tau_{2}=\tau$ in the linearized equation (3), we get
$D^{\alpha}x(t)=(k-\gamma)x(t)-ke^{-\gamma\tau}x(t-\tau),\quad 0<\alpha\leq 1.$ (5)
The above equation is of the general form
$D^{\alpha}x(t)=ax(t)+b(\tau)x(t-\tau),$ (6)
where $a=k-\gamma$ and $b(\tau)=-ke^{-\gamma\tau}$ depends on the delay. We utilize the existing stability results (Theorem 2.1) for the general form (6) for any fixed value of $\tau$ to determine regions of delay-independent stability and instability.
The following theorem summarizes the stability regions of the equilibrium $x_{*}=0$ of (2) for this case in the $k-\gamma$ plane.
Theorem 4.1.
Consider the equation (5) with parameters $k,\gamma\in\mathbb{R}$ and $\tau\geq 0$. Then the equilibrium $x_{*}=0$ has the following behaviors (cf. Figure  4):
  1. (a)
    Delay-Independent instability. If either of the following holds:
    1. (i)
      $\gamma<k<0$, or
    2. (ii)
      $k<\gamma<0$,
    i.e., $(k,\gamma)$ lies in the third quadrant, then the equilibrium is unstable for all $\tau\geq 0$.
  2. (b)
    Delay-Independent stability. The equilibrium is asymptotically stable for all $\tau\geq 0$ if either:
    1. (i)
      $\gamma>2k>0$, or
    2. (ii)
      $\gamma>0$ and $k<0$ i.e. the second quadrant of $k-\gamma$ plane.
  3. (c)
    Delay-Dependent stability.
    1. (i)
      Let $0<\gamma<2k$. Define
      $\tau_{2*}=\frac{-1}{\gamma}\log\left(\frac{|k-\gamma|}{k}\right),\;\;k\neq\gamma\;\;\text{and}$
      $\tau_{1}^{*}(\tau)=\frac{\arccos\left(\dfrac{(k-\gamma)\cos\left(\frac{\alpha\pi}{2}\right)+\sqrt{(-ke^{-\gamma\tau})^{2}-(k-\gamma)^{2}\sin^{2}\left(\frac{\alpha\pi}{2}\right)}\cdot\cos\left(\frac{\alpha\pi}{2}\right)-(k-\gamma)}{-ke^{-\gamma\tau}}\right)}{\left((k-\gamma)\cos\left(\frac{\alpha\pi}{2}\right)+\sqrt{(-ke^{-\gamma\tau})^{2}-(k-\gamma)^{2}\sin^{2}\left(\frac{\alpha\pi}{2}\right)}\right)^{1/\alpha}}$
      Then the stability of the equilibrium is given by:
      • For $\tau<\tau_{2}^{*}$, the system is:
        • Stable if $\tau<\tau_{1}^{*}(\tau)$,
        • Unstable if $\tau>\tau_{1}^{*}(\tau)$.
      • For $\tau>\tau_{2}^{*}$, the system is:
        • Unstable if $0<\gamma<k$,
        • Stable if $k<\gamma<2k$.
    2. (ii)
      Let $k>0,\gamma<0\;\text{i.e. }$ fourth quadrant of the $k-\gamma$ plane. Then, the stability is given by:
      • For $\tau<\tau_{2}^{*}$, equilibrium point is unstable
      • For $\tau>\tau_{2}^{*}$, the equilibrium is
        • stable if $\tau<\tau_{1}^{*}(\tau)$,
        • unstable if $\tau>\tau_{1}^{*}(\tau)$.
Proof.
(a) In both the subcases a(i) and a(ii), $k<0$ and $\gamma<0$. Therefore, $e^{-\gamma\tau}>1$ and $b(\tau)=-ke^{-\gamma\tau}>0$, $\forall\;\tau>0$.
Case (i): $\gamma<k<0\Rightarrow a=k-\gamma>0$.
$\therefore\quad b(\tau)>0>-a\quad\Rightarrow\quad b(\tau)\in(-a,\infty).$
[…]
Case (ii): $k<\gamma<0\Rightarrow a=k-\gamma<0$. Now, define the function $f(\tau):=b(\tau)+a$
$\text{i.e.}\;\;f(\tau):=-ke^{-\gamma\tau}+(k-\gamma).$
We have,
$f(0)=-k+k-\gamma=-\gamma>0.$
As $\tau\rightarrow\infty,$ $f(\tau)\rightarrow\infty$. Differentiating $f(\tau),$ we get
$f^{\prime}(\tau)=k\gamma e^{-\gamma\tau}>0\quad(\text{since }k<\gamma<0).$
Thus, $f(\tau)>0$ for all $\tau$, i.e.,
$b(\tau)>-a.$
Hence, $b(\tau)\in(-a,\infty)$ and again by Theorem 2.1 the system is
unstable for all $\tau>0$.
(b) We show that $a<0$ and $|b(\tau)|<-a$ in both cases b(i) and b(ii), which will ensure the global stability via Part (3) of Theorem 2.1.
Case (i): $\gamma>2k>0\Rightarrow k>0\Rightarrow a=k-\gamma<0$.
Also, $b(\tau)=-ke^{-\gamma\tau}<0\Rightarrow|b(\tau)|=ke^{-\gamma\tau}<k$.
Since $\gamma>2k\Rightarrow-a=\gamma-k>k$, we conclude:
$|b(\tau)|<-a\quad\text{for all }\tau>0.$
Thus, the $x_{*}=0$ is an asymptotically stable equilibrium point of eq.(5).
Case (ii): $\gamma>0$, $k<0\Rightarrow a=k-\gamma<0$, and $b(\tau)=-ke^{-\gamma\tau}>0$.
Thus,
$|b(\tau)|=-ke^{-\gamma\tau}\leq-k<\gamma-k=-a.$
Thus, $|b(\tau)|<-a$, proving global asymptotic stability.
Case (i):
Assume that $0<\gamma<2k.$ Hence, $2k-\gamma>0\;\;\text{i.e.}\;\;k-\gamma>-k.$
$\therefore\;\;a>b(0).\;\;\;\;\;\qquad\qquad$ (*)
At $\tau=0$, equation (5) becomes
$D^{\alpha}x(t)=-\gamma x(t).$
Since $\gamma>0$, equilibrium is asymptotically stable at $\tau=0$. We analyze the sign of $a$ and the condition $b(\tau)<-|a|$.
  • For $a<0,$ we have a=|a|.a=-|a|. Therefore, By (*4), b(0)<|a|.b(0)<-|a|.
  • For $a=k-\gamma>0,$ we have $k>\gamma>0,$
    $\implies-k<-k+\gamma$
    $\implies b(0)<-a=-|a|$
    Thus, we have, $b(0)<-|a|,\;\forall\;a\in\mathbb{R}$, in this case, i.e. the point $(a,b(0))$ lies in SSR region in the Figure 1.
So, as $\tau$ increases, $b(\tau)=-ke^{-\gamma\tau}<0$ will increase and reach to 0. For sufficiently small values of $\tau$ and for $a<0,$ the point $(a,b(\tau))$ will remain in the region SSR of Figure 1 and by Theorem 2.1, Case 1, we get a critical value $\tau_{cr}$. Substituting $a=k-\gamma\;\text{and}\;b=b(\tau)=-ke^{-\gamma\tau}$ in the expression of $\tau_{cr}$, we can easily obtain the critical value $\tau_{1*}(\tau)$ defined in the statement of this Theorem. As we keep on increasing $\tau,$ the point $(a,b(\tau))$ will cross the boundary $b(\tau)=a$ at some critical value $\tau_{2*}$ of $\tau$ (see Figure 2).
Figure 2: Analysis for the region $0<\gamma<2k$
Hence, for $\tau<\tau_{2*},$
  • equilibrium is stable if $\tau<\tau_{1}^{*}(\tau)$ and
  • unstable if $\tau>\tau_{1}^{*}(\tau)$ (by Theorem 2.1, case 1).
Similarly, for $a>0$, the point $(a,b(0))$ belongs to SSR. As $\tau$ increases, the point $(a,b(\tau))$ crosses the boundary b=ab=-a in Figure 2 at $\tau=\tau_{2*}$ and enters in the unstable region.
So, for $a>0,$ $\tau_{2}^{*}$ will be given by setting $b(\tau_{2}*)=-a$ i.e.
$-ke^{-\gamma\tau_{2}*}=\gamma-k$
$\Rightarrow\tau_{2}*=\frac{-1}{\gamma}\log\left(\frac{k-\gamma}{k}\right).$
For $a<0,$ $\tau_{2}^{*}$ will be given by setting $b(\tau_{2}*)=a.$ Hence,
$\tau_{2}*=\frac{-1}{\gamma}\log\left(\frac{\gamma-k}{k}\right).$
Therefore, we have $\tau_{2*}=\frac{-1}{\gamma}\log\left(\frac{|k-\gamma|}{k}\right),\;\;k\neq\gamma.$
Figure 3: Region: $0<\gamma<2k$
For $\tau>\tau_{2}^{*}$, it will enter in stable region if $a<0$ and unstable region if $a>0$ (see Figure 2).
Now, in the region $0<\gamma<2k,$ $a<0$ means $k-\gamma<0$ i.e. $k<\gamma<2k$ and $a>0$ means $k>\gamma$ i.e $0<\gamma<k.$
Thus, we have, for $\tau>\tau_{2*},$ the equilibrium point is unstable if $0<\gamma<k$ and is asymptotically stable if $k<\gamma<2k$ (see Figure 3).
Case (ii): $k>0$ and $\gamma<0$
At $\tau=0$, equation (2) reduces to
$D^{\alpha}x(t)=-\gamma x(t).$
Since $\gamma<0$, the equilibrium is unstable when $\tau=0$. Now, we have $a=k-\gamma>0$ and $b(\tau)=-ke^{-\gamma\tau}<0$. Thus, initially, the point $(a,b(\tau))$ lies in the unstable region in Figure 1, as $k+\gamma<-k$, (in this case), i.e. a<b(0)-a<b(0).
As $\tau$ increases, the value of $b(\tau)$ decreases since its derivative is $b^{\prime}(\tau)=k\gamma e^{-\gamma\tau}<0$. Eventually, $b(\tau)$ crosses the boundary b=ab=-a and enters the single stable region (SSR) (ref. Figure 1). The transition occurs at a critical value $\tau_{2}^{*}$ determined by the expression.
$\tau_{2}^{*}=\frac{-1}{\gamma}\log\left(\frac{k-\gamma}{k}\right).$
For all $\tau<\tau_{2}^{*}$, the equilibrium remains unstable since the point $(a,b(\tau))$ lies entirely within the unstable region. However, once $\tau>\tau_{2}^{*}$, the system enters the delay-dependent SSR region where stability changes with $\tau$.
Inside this region, another critical value $\tau_{1}^{*}(\tau)$, defined by using $\tau_{cr}$ in Theorem 2.1, appears, governing the transition from stability to instability. Substituting the values $a=k-\gamma$ and $b=-ke^{-\gamma\tau}$ in $\tau_{cr},$ we obtain the expression for $\tau_{1*}(\tau).$
Hence, for $\tau>\tau_{2}^{*}$, the equilibrium is stable if $\tau<\tau_{1}^{*}(\tau)$ and becomes unstable again if $\tau>\tau_{1}^{*}(\tau)$. We can expect the switch Unstable-Stable-Unstable in some cases. ∎
Figure 4: Stability regions for Case 1
4.1 Illustrative Examples for Case 1
Consider FDDE $D^{\alpha}x(t)=(k-\gamma)x(t)+ke^{-\gamma\tau}x(t-\tau).$
Example 4.1 Suppose $\alpha=0.4,\;k=-0.1,\;\gamma=-0.2,\;\tau=1.5$. Here, $k,\gamma<0$. By Theorem 4.1(a), the equilibrium is delay-independently unstable (cf. Figure 5).
Figure 5: $\alpha=0.4,\;k=-0.1,\;\gamma=-0.2,\;\tau=1.5$
Example 4.2 Now, consider $\alpha=0.6,\;k=4,\;\gamma=10,\;\tau=0.8$. Here $\gamma>2k>0$. By Theorem 4.1(b)(i), the equilibrium is delay-independently stable. Hence, Numerical solutions converge to zero (see Figure 6).
Figure 6: $\alpha=0.6,\;k=4,\;\gamma=10,\;\tau=0.8$
Example 4.3 Take $\alpha=0.5,\;k=-5,\;\gamma=3,\;\tau=0.5$. Here $k<0,\;\gamma>0$. Theorem 4.1(b)(ii) ensures global stability, which is verified by numerical simulations (cf. Figure 7).
Figure 7: $\alpha=0.5,\;k=-5,\;\gamma=3,\;\tau=0.5$
Example 4.4 For $\alpha=0.4,\;k=1.5,\;\gamma=1.8,\;\tau=1.2$. Here, $0<\gamma<2k$ and $\tau_{2*}=0.8941$. Since, $\tau>\tau_{2*}$, Theorem 4.1(c) predicts stability (see Figure 8).
Figure 8: $\alpha=0.4,\;k=1.5,\;\gamma=1.8,\;\tau=1.2$
Now, suppose $\tau=0.3<\tau_{2*}$, so we find $\tau_{1*}(\tau)=9.2161$. Again, $\tau<\tau_{1*}$, so the equilibrium is stable.
Figure 9: $\alpha=0.4,\;k=1.5,\;\gamma=1.8,\;\tau=0.3$
Example 4.5 Take, $\alpha=0.7,\;k=3.4,\;\gamma=2.2,\;\tau=0.6$. Here, again $0<\gamma<2k$ and $\tau_{2*}=0.4733$. Here, $\tau>\tau_{2*}$, which implies instability. Numerical Simulations verify this (cf. Figure (10).
Figure 10: $\alpha=0.7,\;k=3.4,\;\gamma=2.2,\;\tau=0.6$
Now, let $\tau=0.2<\tau_{2*}$, so we find $\tau_{1*}(\tau)=0.4244$. Again, $\tau<\tau_{1*}$, so the equilibrium is stable.
Figure 11: $\alpha=0.7,\;k=3.4,\;\gamma=2.2,\;\tau=0.2$
Example 4.6 For $\alpha=0.8,\;k=3,\;\gamma=-5,\;\tau=0.08$. Here $k>0,\;\gamma<0$. So, we find $\tau_{2}^{*}=0.1962$. Since $\tau<\tau_{2}^{*}$, the system is unstable. Numerical experiments confirm this (see Figure 12).
Figure 12: $\alpha=0.8,\;k=3,\;\gamma=-5,\;\tau=0.08$
Now, let $\tau=0.35>\tau_{2}^{*}$; $\tau_{1*}(\tau)=0.038614$. Since, $\tau>\tau_{1*}$ implies an unstable equilibrium point (Figure 13).
Figure 13: $\alpha=0.8,\;k=3,\;\gamma=-5,\;\tau=0.35$
Note: By testing a range of $\tau$ values, we verify that stability is independent of delay; however, for the sake of conciseness, we only show a few representative examples here.
5 Case 2: $\tau_{1}=1$
In this case, both delays are retained. Substituting $\tau_{1}=1$ and $\tau_{2}=\tau$ in eq.(3), we obtain
$D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-\tau-1),\quad 0<\alpha\leq 1.$ (7)
The corresponding characteristic equation is given by
$\lambda^{\alpha}=-\gamma+ke^{-\lambda}-ke^{-\gamma\tau}e^{-(\tau+1)\lambda}.$ (8)
We now present some delay-independent results summarizing stability behaviors for this case.
Theorem 5.1.
Consider equation (7) with parameters $k,\gamma\in\mathbb{R}$ and $\tau\geq 0$. Then the equilibrium point $x_{*}=0$ of eq.(2) exhibits the following stability behaviors (see Figure 15):
  1. (a)
    Delay-Independent Stability: The equilibrium is asymptotically stable for all $\tau\geq 0$ if either
    1. (i)
      $\gamma>2k>0$, or
    2. (ii)
      $\gamma>-2k>0$ .
  2. (b)
    Delay-Independent Instability:
    1. (i)
      If $\gamma<0$ and $k<0$, i.e. $(k,\gamma)$ lies in the third quadrant of the $k-\gamma$ plane, then the equilibrium is unstable for all $\tau\geq 0$.
    2. (ii)
      For $\alpha=1,$ the equilibrium is unstable in the fourth quadrant.
Proof.
(a) Consider first $\tau=0$. The eq.(7) reduces to:
$D^{\alpha}x(t)=-\gamma x(t),$
with characteristic equation $\lambda^{\alpha}=-\gamma$. If $\gamma>0$, then $\gamma<0$, so $\arg(-\gamma)=\pi>\frac{\alpha\pi}{2}$, and the equilibrium is asymptotically stable at $\tau=0$.
Now, assume $\tau>0$, and stability will change if a root of the characteristic equation crosses the imaginary axis. Suppose a purely imaginary root $\lambda=iv$ of eq.(8) exists with $v>0$. Substituting into the characteristic equation, we get
$(iv)^{\alpha}=-\gamma+ke^{-iv}\left(1-e^{-(iv+\gamma)\tau}\right).$
Using polar form, $(iv)^{\alpha}=v^{\alpha}e^{i\frac{\alpha\pi}{2}}$, and separating into real and imaginary parts yields
$v^{\alpha}\cos\left(\frac{\alpha\pi}{2}\right)$ $=-\gamma+k\cos v-ke^{-\gamma\tau}\cos(v+v\tau),$ (9)
$v^{\alpha}\sin\left(\frac{\alpha\pi}{2}\right)$ $=-k\sin v+ke^{-\gamma\tau}\sin(v+v\tau).$ (10)
For the system to be unstable, there must exist a root satisfying both eq.(9) and (10).
Case (i) Suppose $\gamma>2k>0,\;k>0$. Then using bounds $\cos v\leq 1$ and $\cos(v+v\tau)\leq 1$, it can be estimated that
$k\cos v-ke^{-\gamma\tau}\cos(v+v\tau)\leq k+ke^{-\gamma\tau},$
so the RHS of (9) satisfies:
$\text{RHS}\leq-\gamma+k+ke^{-\gamma\tau}.$
Since $\tau>0$ and $\gamma>0$, we know that $e^{-\gamma\tau}<1$. Therefore,
$\gamma+k+ke^{-\gamma\tau}<-\gamma+k+k=-\gamma+2k.$
Finally, under the assumption $\gamma>2k$, it follows that
$\gamma+k+ke^{-\gamma\tau}<-\gamma+2k<0,$
which shows that the right-hand side is strictly negative.
Meanwhile, the LHS is non-negative because $v^{\alpha}>0$ and $\cos(\alpha\pi/2)\geq 0$. This gives a contradiction, so no imaginary root exists, and the system remains stable $\forall\;\tau>0$.
Case (ii) Here, $\gamma>-2k>0,\;k<0$. A similar argument as in case (i) will be used here. The bounds $\cos v\geq-1$ and $\cos(v+v\tau)\leq 1$ implies
$k\cos v-ke^{-\gamma\tau}\cos(v+v\tau)\leq-k-ke^{-\gamma\tau}.$
Hence, RHS of (9) satisfies:
$\text{RHS}\leq-\gamma-k-ke^{-\gamma\tau}.$
Again, since $\gamma>-2k$, we get:
$\gamma-k-ke^{-\gamma\tau}<-\gamma-2k<0.$
Because the LHS of eq.(9) is non-negative. This gives a contradiction, so no imaginary root exists, and the system will remain stable in this region too.
Case (i) Assume $\gamma<0$ and $k<0$. At $\tau=0$, we have $\lambda^{\alpha}=-\gamma>0$, so $\lambda$ has a positive real root, implying instability. Now for $\tau>0$, define
$F(\lambda):=\lambda^{\alpha}+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda},\text{the characteristic function.}$ (11)
At $\lambda=0$:
$F(0)=\gamma-k+ke^{-\gamma\tau}.$
$\because\;\;\gamma<0,\tau>0$,
$\qquad\;\;e^{-\gamma\tau}>1$
$\implies ke^{-\gamma\tau}<k\qquad(\because k<0)$
$\implies ke^{-\gamma\tau}-k<0$
$\implies\gamma-k+ke^{-\gamma\tau}<\gamma<0$
$\implies F(0)<0$
Also, $\lim_{\lambda\to\infty}F(\lambda)=\infty$. By the Intermediate Value Theorem, there exists $\lambda_{*}>0$ such that $F(\lambda_{*})=0$, so a positive real root $\lambda_{*}$ exists, and the equilibrium is unstable for all $\tau\geq 0$.
Case (ii) For $\alpha=1,$ the characteristic function (11) becomes
$F(\lambda):=\lambda+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda}.$
Now, for the fourth quadrant, $\gamma<0$, hence, $\gamma>0$. Also,
$F(-\gamma):=0.$
Since the characteristic function has a positive real root, hence equilibrium is unstable in this case. ∎
In the following theorem, we present a delay-dependent result for this case.
Theorem 5.2.
In the fourth quadrant of the $k\!-\!\gamma$ plane (so $k>0$ and $\gamma<0$), fix the parameters $\alpha>0$ and $\tau>0$. There exists a critical value
$k_{*}\;=\;\frac{(1+\tau)^{1+\tfrac{1}{\tau}}}{\tau e^{\gamma}}\left[\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma\right],$ (12)
such that for every $k>k_{*}$ the characteristic equation
$F(\lambda)=\lambda^{\alpha}+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda}$
has a root with positive real part; hence, the equilibrium is unstable for all such $k$.
Proof.
Consider
$F(\lambda)=L(\lambda)-R(\lambda),$
where $\quad L(\lambda)=\lambda^{\alpha}+\gamma\quad\text{and}\quad R(\lambda)=ke^{-\lambda}\big(1-e^{-(\lambda+\gamma)\tau}\big).\\$ Differentiating $R(\lambda)$ gives
$R^{\prime}(\lambda)=ke^{-\lambda}\Big((1+\tau)e^{-(\lambda+\gamma)\tau}-1\Big).$
So, equating $R^{\prime}$ to 0, we get the unique critical point at
$\lambda_{*}=\frac{\ln(1+\tau)}{\tau}-\gamma>0.$
Differentiating $R(\lambda)$ twice gives
$R^{\prime\prime}(\lambda)=ke^{-\lambda}(1-e^{(-\gamma-\lambda)\tau}(1+\tau)^{2}).$
Substituting the critical value $\lambda_{*}$ in $R^{\prime\prime}(\lambda)$ and simplifying, we get,
$R^{\prime\prime}(\lambda_{*})=-k\,\tau\,e^{\gamma}(1+\tau)^{-1/\tau}<0.$
So, $R(\lambda)$ has a local maximum at $\lambda_{*}$. At this point, we compare $L(\lambda_{*})$ and $R(\lambda_{*})$ to get the critical value of k, say $k_{*}$. At $k=k_{*},$ $L(\lambda_{*})=R(\lambda_{*})$ and hence, $\lambda_{*}$ is a positive root of $F$.
$\therefore\qquad\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma=\frac{k_{*}\tau e^{\gamma}}{(1+\tau)^{\frac{1}{\tau}+1}}$
$\implies k_{*}=\frac{(1+\tau)^{1+\tfrac{1}{\tau}}}{\tau e^{\gamma}}\left[\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma\right].$
Figure 14: Comparison of $L(\lambda)=\lambda^{\alpha}+\gamma$ and $R(\lambda)$. At $k=k_{*}$ they intersect at $\lambda_{*},$ where $R(\lambda)$ attains its maximum. For $k>k_{*},$ $R(\lambda)$ exceeds $L(\lambda)$ near $\lambda_{*}$, ensuring a positive root of $F(\lambda)$ and instability.
Finally, since $R(\lambda)$ has a unique maximum, if $k>k_{*}$ then $\max_{\lambda>0}R(\lambda)=R(\lambda_{*})=\frac{k\tau e^{\gamma}}{(1+\tau)^{\frac{1}{\tau}+1}}>L(\lambda_{*})$. Thus $F(\lambda)=L(\lambda)-R(\lambda)$ becomes negative near $\lambda_{*}$, but for large $\lambda$ one has $F(\lambda)\sim\lambda^{\alpha}>0$. By continuity, there exists some $\lambda>0$ with $F(\lambda)=0$, i.e. a positive root. This root implies instability of the equilibrium whenever $k>k_{*}$. ∎
Figure 15: Stability regions for $\tau_{1}=1$
5.1 Illustrative Examples for Case 2
Consider the FDDE $D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-1-\tau)$.
Example 5.1.1
Take $\alpha=0.9,\;k=2.5,\;\gamma=8.5,\;\tau=0.6$. Since $\gamma>2k>0$, Theorem 5.1(a)(i) guarantees delay-independent stability. Simulations confirm convergence (see Figure (16).
Figure 16: $\alpha=0.9,\;k=2.5,\;\gamma=8.5,\;\tau=0.6$
Example 5.1.2 Now, consider $\alpha=0.7,\;k=-3,\;\gamma=7.5,\;\tau=0.4$. Here $\gamma>-2k>0$. By Theorem 5.1(a)(ii), the system is asymptotically stable for all $\tau$. Numerical solutions confirm this (cf. Figure (17).
Figure 17: $\alpha=0.7,\;k=-3,\;\gamma=7.5,\;\tau=0.4$
Example 5.1.3 For $\alpha=0.8,\;k=-4,\;\gamma=-7,\;\tau=0.08$. Since $k<0,\;\gamma<0$, Theorem 5.1(b) predicts delay-independent instability. The simulations show divergence.
Figure 18: $\alpha=0.8,\;k=-4,\;\gamma=-7,\;\tau=0.08$
Note: By testing a range of $\tau$ values, we verify that stability is independent of delay; however, for the sake of conciseness, we only show a few representative examples here.
Example 5.1.4 For $\alpha=0.8,\;k=6,\;\gamma=-0.4,\;\tau=1.6$. Since $k>0,\;\gamma<0$, according to Theorem 5.2, there exists $k_{*}$ such that equilibrium is unstable for all $k>k_{*}.$ Now, $k_{*}=5.675449$, so, k>kk>k*. Hence, the equilibrium is unstable.
Figure 19: $\alpha=0.8,\;k=6,\;\gamma=-0.4,\;\tau=1.6$
5.2 Illustrative Examples for a Nonlinear Function $g(x)$
Let us consider the nonlinear fractional delay differential equation (2) with
$g(x)=k\sin x,$
which leads to
$D^{\alpha}x(t)=-\gamma x(t)+k\sin\!\bigl(x(t-1)\bigr)-e^{-\gamma\tau_{2}}\,k\sin\!\bigl(x(t-1-\tau)\bigr),\quad 0<\alpha\leq 1.$ (13)
To analyze the stability of the trivial equilibrium $x_{*}=0$, we linearize the system around this point. Since
$g^{\prime}(x)=k\cos x\quad\text{and therefore}\quad g^{\prime}(0)=k,$
the linearized equation near $x_{*}=0$ takes the form
$D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-1-\tau).$
This is exactly of the same type as the linearized equation considered in Section 3. Hence, the stability of the trivial solution is again determined by the conditions given in Theorem 5.1. We now illustrate this with a few examples.
Example 5.2.1. Let $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau=0.55$. Then, equation (13) becomes
$D^{0.35}x(t)=-6x(t)+2.4\sin(t-1)-2.4e^{-3.3}\sin(t-1.55).$
The solutions are stable (cf. Fig. LABEL:3a), in agreement with Theorem 5.1, since here $\gamma>2k>0$.
Figure 20: $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau=0.55$.
Example 5.2.2. Now take $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau=1.6$. In this case,
$D^{0.65}x(t)=-2x(t)-0.8\sin(t-1)+0.8e^{-3.2}\sin(t-2.6).$
The solutions converge to the equilibrium (Fig. 21), again supporting Theorem 5.1, since $\gamma>-2k>0$.
Figure 21: $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau=1.6$.
Example 5.2.3. Finally, consider $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau=0.85$. Equation (13) becomes
$D^{0.88}x(t)=5.7x(t)-3.5\sin(t-1)+3.5e^{5.7\cdot 0.85}\sin(t-1.85).$
Here, the solutions are unstable (Fig. 22), consistent with Theorem 5.1, since $(k,\gamma)$ lies in the third quadrant.
Figure 22: $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau=0.85$.
6 Conclusion
This paper investigated the stability of a fractional-order delay differential equation with a delay-dependent coefficient and two discrete delays. By linearizing the system and analyzing the characteristic equation, we derived both delay-independent and delay-dependent stability conditions. We studied two cases: $\tau_{1}=0$ and $\tau_{1}=1$ in detail. The numerical simulations and stability diagrams supported the theoretical findings.
The results show that the interaction between fractional order, feedback strength, and delay-dependent coefficients has a major influence on system stability. In particular, the presence of such coefficients can cause stability switches and transitions. These insights are useful for modeling biological and control systems where feedback and memory effects are important.
Future studies may extend this work to a deeper bifurcation analysis, nonlinear models with multiple delays, and stabilization methods for unstable states.